### v1

@@ 1 Introduction
Strangeness photoproduction is a crucial tool to understand baryon resonance spectra and the underlying degrees of freedom afforded in the low energy, non-perturbative regime of QCD. $K\Lambda^{*}$ and $K\Sigma^{*}$ photoproduction has frequently been used over the last twenty years to probe the third resonance region in attempts to characterise known higher mass $N^{*}$ and $\Delta$ states and search for so called missing resonances which have been predicted by constituent quark models capstick86 ; capstick92 ; capstick94 ; riska01 ; klempt10 ; klempt12 , lattice QCD calculations edwards11 , and Dyson-Schwinger equations of QCD roberts11 but not observed experimentally in $N\pi$ final states capstick00 ; loering01 . $K\Lambda^{*}$ channels are unique in that the isoscalar $\Lambda$ acts as an isospin filter, only permitting intermediate $s$-channel $N^{*}$ resonances, and for the ground state $\Lambda$, the self-analysing weak decay enables easier access to baryon-recoil polarisation observables.
@@
The $\Lambda^{*}$ excitation spectrum however is itself poorly understood compared to the non-strange baryon sector, mainly due to a scarcity of experimental data pdg20 . Very few measurements on $S=-1$ resonances (both $\Lambda^{*}$ and $\Sigma^{*}$ states) have been made over the last 20 years. The exception to this are measurements of the final states $K^{+}\Sigma^{0}\pi^{0}$ and $K^{+}\Sigma^{\mp}\pi^{\pm}$ lu13 ; moriya13 ; moriya13b ; niiyama08 ; lambda1405paper in both electro- and photoproduction, which have mainly been focussed on decays of the $\Lambda(1405)$. Studies of the $\Lambda(1405)$ are largely driven by its exotic nature and it is now considered the archetypal molecular-like hadron in the $uds$ quark sector guo18 . The large mass difference to the spin-orbit partner, the $\Lambda(1520)$ is also hard to explain within a CQM, and given the strong evidence of the molecular nature of the $\Lambda(1405)$, it is natural to question the structure of the $\Lambda(1520)$. Refs. kolomeitsev04 ; sarkar05 ; sarkar05b , for example employed a chiral coupled channel model which described the $\Lambda(1520)$ as a quasi-bound state of $\pi\Sigma(1385)$ and $K\Xi(1530)$.
@@
Previous measurements of $K^{+}\Lambda(1520)$ photoproduction near threshold have been performed at the LEPS kohri10 and CLAS moriya13b ; shrestha21 experiments [^1] There are also earlier measurements made at 11 GeV at the SLAC spectrometer boyarski71 and between 2.8 and 4.8 GeV from the LAMP2 collaboration barber80 ., which have been supported by models employing effective Lagrangians he12 ; nam05 . The model of He and Chen he12 was able to describe the data by including the contact term, $s$-channel, $u$-channel and $t$-channel $K$ and $K^{*}$ exchanges. It was determined that the contact term dominates close to threshold, with $K^{*}$ $t$-channel contributions at higher centre-of-mass energies. Two $N^{*}$ resonances with spins 3/2 and 5/2 were also required with masses close to 2.1 GeV for an adequate reproduction of the data. The LEPS collaboration also reported a bump like structure at a centre-of-mass energy of approximately 2.1 GeV. It was speculated that this may be due to either interference with the $\phi p$ final state or alternatively an additional resonance. The angular distribution however did not support the introduction of a new resonance at this energy.
@@
If the $\Lambda(1520)$ is a candidate for a molecular-like state, it is of interest to determine the dependency that momentum transfer has upon the differential cross section. It would be expected that a loosely bound system breaks up under large momentum transfer, which is observed for example, in coherent reactions off deuterium and light nuclei (see for example, Ref. krusche11 ). Accessing forward $K^{+}$ angles close to threshold minimises the momentum transfer (minimising the Mandelstam variable, $t$) to the $\Lambda(1520)$, and an enhancement or structure may be expected if a meson-baryon type “resonating structure” contributes, for example, a molecular-type configuration. The BGOOD photoproduction experiment technicalpaper at the ELSA facility hillert06 ; hillert17 at the University of Bonn is ideally suited to measure associated strangeness in this kinematic regime. The $K^{+}$ can be identified at forward angles in the Forward Spectrometer, while the hyperon decays almost isotropically, the decay products of which are identified in the central region with the BGO Rugby Ball electromagnetic calorimeter. The BGOOD collaboration is pursuing a programme of strangeness photoproduction measurements under low-momentum transfer conditions lambda1405paper ; klambdapaper ; ksigmapaper which have not been achievable with high statistical precision at other facilities. An example of the importance of measuring this region is the measured forward differential cross section for $K^{+}\Sigma^{0}$, where a cusp-like structure was observed at the $K^{+}\Sigma(1385)$ threshold ksigmapaper ; INPCC22Conf . The measurement of $K^{+}\Lambda(1520)$ photoproduction close to threshold at forward angles is therefore the mandatory next step in order to understand the role of potential molecular-like structure in the strangeness sector.
@@ 2 Experimental setup and analysis procedure
The presented results are from data taken over two periods. The first dataset was taken in 2017 and used a 6 cm long liquid hydrogen target with a 3.2 GeV electron beam over 22 days, resulting in an integrated photon flux of $1.11\times 10^{12}$ over the photon energy range 1650 to 1850 MeV (the region of the presented data). The second dataset taken in 2018 used an 11 cm long liquid hydrogen target with a 2.8 GeV electron beam over 18 days, resulting in an integrated photon flux of $7.28\times 10^{11}$ over the photon energy range 1650 to 1850 MeV. Four different hardware triggers were implemented to ensure the data can be used a variety of measurements (see Refs. technicalpaper ; klambdapaper ), however the results presented here required a single trigger combination, which was an incident photon registered in the photon tagger and a minimum energy deposition of 150 MeV in the BGO Rugby Ball.
$K^{+}$ were identified in the Forward Spectrometer by reconstructing their invariant mass from the momentum and $\beta$ measurements. $K^{+}$ candidates were selected over a $2\sigma$ range of the reconstructed mass. This increased linearly with $K^{+}$ momentum, ranging from 45 MeV/c2 to 73 MeV/c2 at momenta 450 and 1000 GeV/c respectively.
@@
The event yields contained background from hadronic reactions, where either the forward $K^{+}$ was misidentified from the large background of $\pi^{+}$ and $e^{+}$, or other associated strangeness reactions of $K^{+}Y$ and $K^{+}Y\pi$ were selected instead of $K^{+}\Lambda(1520)$. This background can be observed in histograms of the missing mass recoiling from the $K^{+}$ in the final state. Three examples for different centre-of-mass energy ($W$) intervals are shown in Fig. 2 for the two different topologies (a $K^{+}$ candidate with or without an additional $\pi^{0}$ candidate). To extract the $K^{+}\Lambda(1520)$ yields, these were fitted with distributions from simulated associated strangeness channels. This was achieved using the BGOOD GEANT4 geant4 simulation, including all spatial, energy and time resolutions, magnetic fields and hardware efficiencies (see Ref. technicalpaper for details). Event distributions were based on measured data where available or phase distributions, however the small $W$ and $\cos\theta_{\mathrm{CM}}^{K}$ range per interval made differences in distributions negligible. Signal events ($K^{+}\Lambda(1520)$) and background from $K^{+}\Lambda$, $K^{+}\Sigma^{0}$, $K^{+}\Sigma^{0}(1385)$ and $K^{*+}\Lambda$ were simulated. $K^{+}\Lambda(1405)$ was not included due to the mass degeneracy to the $\Sigma(1385)$. From previous studies klambdapaper , it was realised that $\Delta\pi^{+}$ was required as it gave substantial background contribution from $K^{+}$-$\pi^{+}$ misidentification, however this only contributes at lower missing mass than the $\Lambda(1520)$ signal. An additional background distribution from misidentifying either $\pi^{+}$ or $e^{+}$ as $K^{+}$ was determined by repeating the analysis but requiring a negatively charged track (a technique first used in Refs. klambdapaper ; ksigmapaper ). This resulted in $\pi^{-}$ and $e^{-}$ distributions which appeared (by eye) identical to their positively charged counterparts. The measured data extends only to the $K^{*}\Sigma^{0}$ threshold as the missing mass distribution from this background channel was too close to the $K^{+}\Lambda(1520)$ distribution for a reliable separation of signal and background. [^3] Preliminary measurements of $K^{*}Y$ channels at BGOOD have been made and are under analysis for future publications. Once complete, it is the intention to subtract the $K^{*}\Sigma^{0}$ contributions and extend the measured $K^{+}\Lambda(1520)$ $W$ range.
@@ 3 Results and interpretation
The differential cross section for $\cos\theta_{\mathrm{CM}}^{K}$ $>0.9$ is shown in Fig. 5 (filled black circles). The data provide the first measurements with high precision close to threshold, achieving a resolution in $W$ of approximately 10 MeV which is 2.5 times finer than previous LEPS data kohri10 with similar statistical precision. Superimposed is the effective Lagrangian model of He and Chen he12 which was fitted to the LEPS data. There is a good agreement to the model calculation, with a reduced $\chi^{2}$ of 1.96 when only including the statistical errors, however when combining the systematic and statistical errors in quadrature, a reduced $\chi^{2}$ of 1.04 is achieved. The data support the case for the contact term to dominate the cross section close to threshold.
@@
Figure 5: The differential cross section for $\cos\theta_{\mathrm{CM}}^{K}$$>0.9$ for the combined BGOOD data set requiring only a forward $K^{+}$ (filled black circles with systematic errors indicated as the filled grey area on the abscissa). Previous LEPS kohri10 and CLAS data moriya13b are shown as filled red squares and filled blue triangles respectively. For all data, vertical bars indicate statistical error and horizontal bars indicate the range of the data point. The CLAS data covers a $\cos\theta_{\mathrm{CM}}^{K}$ range of 0.8 to 0.9. The results from the effective Lagrangian model of He and Chen he12 are superimposed and labelled inset.
@@
The $\cos\theta_{\mathrm{CM}}^{K}$ resolution of approximately 0.02 permits binning in finer $\cos\theta_{\mathrm{CM}}^{K}$ intervals. Fig. 6 shows the same dataset in $\cos\theta_{\mathrm{CM}}^{K}$ intervals of 0.033. No significant change is observed across this $\cos\theta_{\mathrm{CM}}^{K}$ range, which is in contrast, for example, to the $K^{+}\Sigma^{0}$ differential cross section which exhibits a cusp-like structure becoming more prominent in the most forward interval ksigmapaper . This fine binning in $\cos\theta_{\mathrm{CM}}^{K}$ provides an accurate means to extrapolate the data to the minimum possible momentum transfer to the $\Lambda(1520)$. This is defined in terms of the Mandelstam variable, $t=(p_{\gamma}-p_{K})^{2}$, where $p_{\gamma}$ and $p_{K}$ are the four-momenta of the photon beam and $K^{+}$ respectively. To account for the distribution of $t$ within each two dimensional $W$ and $\cos\theta_{\mathrm{CM}}^{K}$ interval, a generated distribution was used based on the measured differential cross section. For each interval in $W$ and $\cos\theta_{\mathrm{CM}}^{K}$, the mean average value of $t$ was used as the central value, and the width was determined as $\sqrt{12}$ RMS. A fixed mass of the $\Lambda(1520)$ was assumed for the calculation in order to remove long tails to very high and very small $t$. For each $W$ interval, the minimum momentum transfer, $t_{\mathrm{min}}$ was calculated, which corresponds to $\cos\theta_{\mathrm{CM}}^{K}$= 1. Fig. 7 shows examples of the differential cross section with respect to $|t-t_{\mathrm{min}}|$ for three different $W$ intervals.
@@
Figure 7: The differential cross section with respect versus $|t-t_{\mathrm{min}}|$ for three example intervals of $W$ (only the statistical error shown). Equation 1 has been fitted separately to each $W$ interval, indicated by the colour coded lines.
@@
Fig. 8(a) shows the extrapolated differential cross section at $t_{\mathrm{min}}$. The slope parameter, $S$ is shown in Fig. 8(b). Within limited statistical precision it appears to remain flat. This is consistent with the contact term dominating near threshold, as was proposed by the effective Lagrangian model of He and Chen he12 . If $K$ or $K^{*}$ $t$-channel exchange was to dominant for example, the differential cross section would be expected to be more forward peaked, resulting in a negative value for $S$.

### v2

@@ 1 Introduction
Strangeness photoproduction is a crucial tool to understand baryon resonance spectra and the underlying degrees of freedom afforded in the low energy, non-perturbative regime of QCD. $K\Lambda^{*}$ and $K\Sigma^{*}$ photoproduction has frequently been used to probe the third resonance region in attempts to characterise known higher mass $N^{*}$ and $\Delta$ states and search for so called missing resonances which have been predicted by constituent quark models capstick86 ; capstick92 ; capstick94 ; riska01 ; klempt10 ; klempt12 , lattice QCD calculations edwards11 , and Dyson-Schwinger equations of QCD roberts11 but not observed experimentally in $N\pi$ final states capstick00 ; loering01 . $K\Lambda^{*}$ channels are unique in that the isoscalar $\Lambda$ acts as an isospin filter, permitting only intermediate $s$-channel $N^{*}$ resonances, and in the case of the ground state $\Lambda$, the self-analysing weak decay enables easier access to baryon-recoil polarisation observables.
@@
The $\Lambda^{*}$ excitation spectrum however is itself poorly understood compared to the non-strange baryon sector, mainly due to a scarcity of experimental data pdg20 . Very few measurements on $S=-1$ resonances (both $\Lambda^{*}$ and $\Sigma^{*}$ states) have been made over the last 20 years. The exception to this are measurements of the final states $K^{+}\Sigma^{0}\pi^{0}$ and $K^{+}\Sigma^{\mp}\pi^{\pm}$ lu13 ; moriya13 ; moriya13b ; niiyama08 ; lambda1405paper in both electro- and photoproduction, which have mainly focussed on decays of the $\Lambda(1405)$. Studies of the $\Lambda(1405)$ are largely driven by its exotic nature and it is now considered the archetypal molecular-like hadron in the $uds$ quark sector guo18 . The large mass difference to the spin-orbit partner, the $\Lambda(1520)$ is also hard to explain within a CQM, and given the strong evidence of the molecular nature of the $\Lambda(1405)$, it is natural to question the structure of the $\Lambda(1520)$. Refs. kolomeitsev04 ; sarkar05 ; sarkar05b , for example employed a chiral coupled channel model which described the $\Lambda(1520)$ as a quasi-bound state of $\pi\Sigma(1385)$ and $K\Xi(1530)$.
@@
Previous measurements of $K^{+}\Lambda(1520)$ photoproduction near threshold were made at the LEPS kohri10 , CLAS moriya13b ; shrestha21 and SAPHIR wieland11 experiments [^1] There are also earlier measurements made at 11 GeV at the SLAC spectrometer boyarski71 and between 2.8 and 4.8 GeV from the LAMP2 collaboration barber80 ., which have been supported by models employing effective Lagrangians he12 ; nam05 ; nam07 ; nam10 ; toki08 ; xie10 ; xie14 ; wie21 and models including additional Regge trajectories to describe higher energies nam10b ; wang14 ; he14 ; yu17 . These analyses generally agree that near threshold the contact term dominates, with $t$-channel $K$ exchange important at higher energies. Contributions from $t$-channel $K^{*}$ exchange, $u$-channel $\Lambda$ exchange and varying $s$-channel $N^{*}$ contributions however vary significantly, largely due to the unknown $\bar{K}^{*}N\Lambda^{*}$ coupling and choice of hadronic form factors. The model of He and Chen he12 for example, was used to determine that the contact term dominates close to threshold, with $K^{*}$ $t$-channel contributions at higher centre-of-mass energies. Two $N^{*}$ resonances with spins 3/2 and 5/2 were also required with masses close to 2.1 GeV for an adequate reproduction of the data. Similarly, the model of Xie, Wang and Nieves supported the contribution of an $N^{*}(2120)$ with $J^{P}=3/2^{-}$ xie10 ; xie14 . This corroborates the results of a coupled channel model of Ramos and Oset ramos13 , which suggested a dynamically generated state at 2 GeV with $J^{P}=3/2^{-}$ plays a dominant role in $K^{0}\Sigma$ photoproduction ewald12 ; k0paper . The analyses by Wei et al. wie21 employed the first effective Lagrangian model to make a global fit to differential cross section and beam asymmetry data, with the strong motivation that previous effective Lagrangian models were unable to describe LEPS beam asymmetry measurements kohri10 . Two fits were made, either including an $N(2050)5/2^{-}$ or $N(2120)3/2^{-}$ resonance (denoted Fit A and B respectively), both giving satisfactory results. Whilst the contact term and $t$-channel $K$ exchange were dominant in both models, contributions from $K^{*}$ and $u$-channel $\Lambda$ exchange varied markedly.
@@
If the $\Lambda(1520)$ is a candidate for a molecular-like state, it is of interest to determine the dependency that momentum transfer has upon the differential cross section. It would be expected that a loosely bound system breaks up under large momentum transfer, which is observed for example, in coherent reactions off the deuteron and light nuclei (see for example, Ref. krusche11 ). Accessing forward $K^{+}$ angles close to threshold minimises the momentum transfer (minimising the Mandelstam variable, $t$) to the $\Lambda(1520)$, and an enhancement or structure may be expected if a meson-baryon type “resonating structure” contributes, for example, a molecular-type configuration. The BGOOD photoproduction experiment technicalpaper at the ELSA facility hillert06 ; hillert17 at the University of Bonn is ideally suited to measure associated strangeness in this kinematic regime. The $K^{+}$ can be identified at forward angles in the Forward Spectrometer, while the hyperon decays almost isotropically, the decay products of which are identified in the central region with the BGO Rugby Ball electromagnetic calorimeter. The BGOOD collaboration is pursuing a programme of strangeness photoproduction measurements under low-momentum transfer conditions lambda1405paper ; klambdapaper ; ksigmapaper which have not been achievable with high statistical precision at other facilities. An example of the importance of studying this region includes the forward differential cross section for $K^{+}\Sigma^{0}$, where a cusp-like structure was observed at the $K^{+}\Sigma(1385)$ threshold ksigmapaper ; INPCC22Conf . The measurement of $K^{+}\Lambda(1520)$ photoproduction close to threshold at forward angles is therefore the mandatory next step in order to understand the role of potential molecular-like structure in the strangeness sector.
@@ 2 Experimental setup and analysis procedure
The presented results are from data taken over two periods. The first dataset was taken in 2017 and used a 6 cm long liquid hydrogen target with a 3.2 GeV electron beam over 22 days, resulting in an integrated photon flux of $1.11\times 10^{12}$ over the photon energy range 1650 to 1850 MeV (the region of the presented data). The second dataset taken in 2018 used an 11 cm long liquid hydrogen target with a 2.8 GeV electron beam over 18 days, resulting in an integrated photon flux of $7.28\times 10^{11}$ over the same photon energy range. Four different hardware triggers were implemented to ensure the data can be used for a variety of measurements (see Refs. technicalpaper ; klambdapaper ), however the results presented here required a single trigger combination, which was an incident photon registered in the photon tagger and a minimum energy deposition of 150 MeV in the BGO Rugby Ball.
$K^{+}$ were identified in the Forward Spectrometer by reconstructing their invariant mass from momentum and $\beta$ measurements. $K^{+}$ candidates were selected over a $2\sigma$ range of the reconstructed mass. This increased linearly with $K^{+}$ momentum, ranging from 45 MeV/c2 to 73 MeV/c2 at momenta 450 and 1000 MeV/c respectively.
@@
The event yields contained background from hadronic reactions, where either the forward $K^{+}$ was misidentified from $\pi^{+}$ or $e^{+}$, or other associated strangeness reactions of $K^{+}Y$ and $K^{+}Y\pi$. This background can be observed in histograms of the missing mass recoiling from the $K^{+}$ in the final state, examples of which are shown in Fig. 2 for the two different topologies (a $K^{+}$ candidate with or without an additional $\pi^{0}$ candidate). To extract the $K^{+}\Lambda(1520)$ yields, these were fitted with distributions from simulated associated strangeness channels. This was achieved using the BGOOD GEANT4 geant4 simulation, including all spatial, energy and time resolutions, magnetic fields and hardware efficiencies (see Ref. technicalpaper for details). Event distributions were based on measured data where available or phase distributions, however the small $W$ and $\cos\theta_{\mathrm{CM}}^{K}$ range per interval made differences in distributions negligible. Signal events ($K^{+}\Lambda(1520)$) and background from $K^{+}\Lambda$, $K^{+}\Sigma^{0}$, $K^{+}\Sigma^{0}(1385)$ and $K^{*+}\Lambda$ were simulated. $K^{+}\Lambda(1405)$ was not included due to the mass degeneracy to the $\Sigma(1385)$. From previous studies klambdapaper , it was realised that $\Delta\pi^{+}$ was required as it gave substantial background contribution from $K^{+}$-$\pi^{+}$ misidentification, however this only contributes at lower missing mass than the $K^{+}\Lambda(1520)$ signal. An additional background distribution from misidentifying either $\pi^{+}$ or $e^{+}$ as $K^{+}$ was determined by repeating the analysis but requiring a negatively charged track (a technique first used in Refs. klambdapaper ; ksigmapaper ). This resulted in $\pi^{-}$ and $e^{-}$ distributions which appeared (by eye) identical to their positively charged counterparts. The measured data extends only to the $K^{*}\Sigma^{0}$ threshold as the missing mass distribution from this background channel was too close to the $K^{+}\Lambda(1520)$ distribution for a reliable separation of signal and background. [^3] Preliminary measurements of $K^{*}Y$ channels at BGOOD have been made and are under analysis for future publications. Once complete, it is the intention to subtract the $K^{*}\Sigma^{0}$ contributions and extend the measured $K^{+}\Lambda(1520)$ $W$ range.
@@ 3 Results and interpretation
The differential cross section for $\cos\theta_{\mathrm{CM}}^{K}$ $>0.9$ is shown in Fig. 5 (filled black circles). The data provide the first measurements with high precision close to threshold, achieving a resolution in $W$ of approximately 10 MeV which is 2.5 times finer than previous LEPS data kohri10 with similar statistical precision. Superimposed are the effective Lagrangian models of He and Chen he12 , Xie and Nieves xie10 ; xie14 and Wei et. al wie21 which were fitted to previous LEPS and CLAS data and give good agreement to the BGOOD data. The He and Chen model, for example has a reduced $\chi^{2}$ of 1.96 when only including the statistical errors, however when combining the systematic and statistical errors in quadrature, a reduced $\chi^{2}$ of 1.04 is achieved. The model of Wie et al. used a global fit to differential cross section data and beam asymmetry measurements for both Fits A and B (including either $N(2060)5/2^{-}$ or $N(2120)3/2^{-}$ respectively), with both giving a reasonable agreement. In all cases, the BGOOD data support the case for the contact term to dominate the cross section close to threshold, and it is anticipated that the improved statistical precision will further constrain effective Lagrangian fit parameters in future analyses.
@@
Figure 5: The differential cross section for $\cos\theta_{\mathrm{CM}}^{K}$$>0.9$ for the combined BGOOD data set requiring only a forward $K^{+}$ (filled black circles with systematic errors indicated as the filled grey area on the abscissa). Previous LEPS kohri10 and CLAS data moriya13b (at the more backward $\cos\theta_{\mathrm{CM}}^{K}$ of 0.8 to 0.9) are shown as filled red squares and filled blue triangles respectively. For all data, vertical bars indicate statistical error and horizontal bars indicate the range of the data point. The results from the effective Lagrangian models of Xie and Nieves xie10 , He and Chen he12 and Wei et. al. wie21 are superimposed and labelled inset.
@@
The $\cos\theta_{\mathrm{CM}}^{K}$ resolution of approximately 0.02 permits binning in finer $\cos\theta_{\mathrm{CM}}^{K}$ intervals. Fig. 6 shows the same dataset in $\cos\theta_{\mathrm{CM}}^{K}$ intervals of 0.033. No significant change is observed across this $\cos\theta_{\mathrm{CM}}^{K}$ range, which is in contrast, for example, to the $K^{+}\Sigma^{0}$ differential cross section which exhibits a cusp-like structure becoming more prominent in the most forward interval ksigmapaper . This fine binning in $\cos\theta_{\mathrm{CM}}^{K}$ provides an accurate means to extrapolate the data to the minimum possible momentum transfer to the $\Lambda(1520)$. This is defined in terms of the Mandelstam variable, $t=(p_{\gamma}-p_{K})^{2}$, where $p_{\gamma}$ and $p_{K}$ are the four-momenta of the photon beam and $K^{+}$ respectively. To account for the distribution of $t$ within each two dimensional $W$ and $\cos\theta_{\mathrm{CM}}^{K}$ interval, a generated distribution was used based on the measured differential cross section. For each interval in $W$ and $\cos\theta_{\mathrm{CM}}^{K}$, the mean average value of $t$ was used as the central value, and the width was determined as $\sqrt{12}$ RMS. A fixed mass of the $\Lambda(1520)$ was assumed for the calculation in order to remove long tails to very high and very small $t$. For each $W$ interval, the minimum momentum transfer, $t_{\mathrm{min}}$ was calculated, corresponding to $\cos\theta_{\mathrm{CM}}^{K}$= 1. Fig. 7 shows examples of the differential cross section with respect to $|t-t_{\mathrm{min}}|$ for three different $W$ intervals.
@@
Figure 7: The differential cross section with respect to $t$ versus $|t-t_{\mathrm{min}}|$ for three example intervals of $W$ (only the statistical error shown). Equation (1) is fitted to each $W$ interval, indicated by the colour coded lines.
@@
Fig. 8(a) shows the extrapolated differential cross section at $t_{\mathrm{min}}$. The slope parameter, $S$ is shown in Fig. 8(b). Within limited statistical precision it appears to remain flat. This is consistent with the contact term dominating near threshold, as was proposed by many effective Lagrangian models he12 ; nam05 ; nam07 ; nam10 ; toki08 ; xie10 ; xie14 . If $K$ or $K^{*}$ $t$-channel exchange was dominant for example, the differential cross section could be expected to be more forward peaked, resulting in a negative value for $S$.